OpenStudy (mopsy510):

The radius of a circle is 1 inch. What is the area of a sector bounded by a 75° arc?

OpenStudy (anonymous):

n/360 (pi)(radius)^2 = 75/360 (pi)(1)^2 = 5pi/24

OpenStudy (anonymous):

5pi/24 square inches